Section 15  Numerical Analysis and Scientific Computing

Section 15: Numerical Analysis Scientific Computing

Invited Lectures

Monday, 24 August  H 1028

16:15 – 17:00 Hisashi Okamoto, RIMS, Kyoto University, Kyoto, Japan Bifurcation of Kolmogorov flows with an emphasis on singular limit (I: p.180) Chair: Rolf Jeltsch

17:15 – 18:00 Gregory Beylkin, University of Colorado at Boulder, Boulder, USA On multiresolution methods in numerical analysis (I: p.178) Chair: Siegfried Prößdorf

Tuesday, 25 August  H 1012

14:00 – 14:45 Percy Alec Deift, Courant Institute, New York University, New York, USA Uniform asymptotics for orthogonal polynomials (I: p.179) Chair: Tom H. Koornwinder

15:00 – 15:45 Jan-Olov Strömberg, Royal Institute of Technology, Stockholm, Sweden Apriori estimates for approximation by wavelets in higher dimensions (I: p.180) Chair: Peter Maaß

Wednesday, 26 August  H 1012

14:00 – 14:45 Lloyd N. Trefethen, Oxford University, Oxford, UK Schwarz–Christoffel Mapping in the Computer Era (I: p.181) Chair: Peter J. Deuflhard

15:00 – 15:45 Bjorn Erik Engquist, UCLA, Los Angeles, USA Wavelet Based Numerical Homogenization (I: p.179) Chair: Ralf Kornhuber

Numerical Analysis and Scientific Computing  Section 15

Short Communications and Poster Sessions

Short Communications  Wednesday, 19 August
Ordinary Differential Equations I  H 2013

Chair: Alexander Ostermann

14:00 – 14:15 Gennady Yu. Kulikov, Ulyanovsk State University, Ulyanovsk, Russia How to control the global error of Runge-Kutta methods (p.305)

14:15 – 14:30 Alexander Ostermann, Universität Innsbruck, Innsbruck, Austria The unstable manifold of an equilibrium point under discretization (p.308)

14:30 – 14:45 George Micula, Babes-Bolyai University, Cluj-Napoca, Romania Polynomial spline functions of even degree approximating the solution of differential equations

Ordinary Differential Equations II  H 2013

Chair: Tatiana Anatolyevna Vasilyeva

15:00 – 15:15 Felix Schwitzer, Universität Innsbruck, Innsbruck, Austria Linearly implicit Runge-Kutta discretizations of semilinear parabolic problems near equilibria

(p.312)

15:15 – 15:30 Sergei Yu. Pilyugin, St.-Petersburg University, St.-Petersburg, Russia Shadowing and perturbation of attractors in discretized parabolic equations (p.310)

15:30 – 15:45 Tatiana Anatolyevna Vasilyeva, Volgograd State University, Volgograd, Russia Twice implicit scheme with accuracy control and it application to chemical kinetics problems

(p.316)

Wavelets I (organized by Andreas Rathsfeld)  H 2033

Chair: Wayne M. Lawton

16:15 – 16:30 Jacques Liandrat, Univ Aix Marseille II, Marseille, France Wavelets, from theory to efficients numerical algorithms (p.306)

16:30 – 16:45 Andreas Rieder, Universität des Saarlandes, Saarbrücken, Germany Embedding and a-priori adaptivity (p.311)

16:45 – 17:00 Pierre Gilles Lemarié-Rieusset, Université d’Evry, Evry, France Some remarks on divergence free vector wavelets (p.306)

Nonlinear Operator Equations  H 2013

Chair: Jose A. Ezquerro

16:15 – 16:30 Jose A. Ezquerro, University of la Rioja, Logrono, Spain A semilocal convergence result for Secant-type methods under Hölder continuity conditions (p.299)

16:30 – 16:45 Anatolii Borisovich Bakushinskii, Inst. for System Analysis, Moscow, Russia Iterative methods for the solution of nonlinear operator equations without property of the regularity

(p.296)

16:45 – 17:00 Victor D. Didenko, Odessa State University, Odessa, Ukraine On Peculiarities of Approximation Methods for Singular Integral Equations with Conjugation

(p.298)

Interpolation  H 2033

Chair: Raquel R. Martinez

17:15 – 17:30 Raquel R. Martinez, Universidad Politécnica de Madrid, Madrid, Spain Composite Methods for Interpolating Functions (p.307)

17:30 – 17:45 Wolfgang Volk, Berlin, Germany Quasi-interpolation with Arbitrarily Often Differentiable Functions (p.316)

17:45 – 18:00 Yurij Kaluch I., Ukrainian Academy Of Public Administration, Kiev, Ukraine Mathematical Problems of Condition Prognosing of Land Slopes and Slides (p.303)

Short Communications  Thursday, 20 August
Approximation I  H 2033

Chair: Jozsef Grof

14:00 – 14:15 Alexis Bacopoulos, National Technical University, New Psychico, Athens, Greece On the link among vectorial minimal projections and best approximations (p.296)

14:15 – 14:30 Igor Fedotov, University of Transkei, Umtata, South Africa Another Approach to Quadrature Methods (p.300)

14:30 – 14:45 Jozsef Grof, Universität Veszprem, Veszprem, Hungary Einige Verallgemeinerungen bzw. Modifizierungen der Szász-Mirakjan-Operatoren (p.301)

Partial Differential Equations I  H 2013

Chair: Vladimir V. Shaidurov

14:00 – 14:15 Christoph Pflaum, Institut für Angewandte Mathematik und Statistik, Würzburg, Germany Robust Multilevel Algorithms for Elliptic Differential Equations (p.309)

14:15 – 14:30 Vladimir V. Shaidurov, Russian Academy of Sciences, Krasnoyarsk, Russia Cascade finite element methods for weakly nonlinear elliptic equations (p.313)

14:30 – 14:45 Urve Kangro, Tartu University, Tartu, Estonia Spectral Convergence of the Interior Source Method with Point Matching (p.303)

Approximation II  H 2033

Chair: Alexei A. Mailybaev

15:00 – 15:15 Alois Heinz, Universität Freiburg, Freiburg, Germany Efficient function approximation by tree-structured neural networks (p.302)

15:15 – 15:30 Peter Kirrinnis, Universität Bonn, Bonn, Germany Fast algorithms for partial fractions (p.305)

15:30 – 15:45 Alexei A. Mailybaev, Lomonosov Moscow State University, Moscow, Russia Transformation of Families of Matrices to Normal Forms (p.306)

Partial Differential Equations II  H 2013

Chair: Jürgen Fuhrmann

15:00 – 15:15 Bodo Erdmann, Konrad-Zuse-Zentrum, Berlin, Germany Mesh Controlling Strategy for Solving Nonlinear Evolution Problems (p.299)

15:15 – 15:30 Christian Wieners, University Stuttgart, Stuttgart, Germany Adaptive multigrid methods for Prandtl-Reuß-plasticity (p.318)

15:30 – 15:45 Jürgen Fuhrmann, Weierstraß Institut, Berlin, Germany Numerically Stable Finite Volume Discretizations for Nonlinear Parabolic Problems (p.300)

Wavelets II (organized by Andreas Rathsfeld)  H 2013

Chair: Jacques Liandrat

16:15 – 16:30 Rudolph A. Lorentz, GMD, St. Augustin, Germany On the nonexistence of compactly supported Sobolev prewavelets for box spline multiresolution analyses (p.306)

16:30 – 16:45 Gunther Schmidt, WIAS, Berlin, Germany Representation of multidimensional potentials by approximate wavelets (p.312)

16:45 – 17:00 Wayne M. Lawton, National University, Singapore, Singapore Energy functions on graphs, wavelets, and multilevel algorithms (p.305)

Differential Algebraic Systems  H 2013

Chair: Ulrich Nowak

17:15 – 17:30 Rainald Ehrig, Konrad-Zuse-Zentrum für Informationstechnik, Berlin, Germany Recent advances in extrapolation methods for large scale differential algebraic problems (p.299)

17:30 – 17:45 Friedrich Grund, Weierstraß Institut, Berlin, Germany Sparse-Matrix-Techniken für Parallelrechner (p.301)

17:45 – 18:00 Tatyana Shtykel, Humboldt–Universität, Berlin, Germany On the criterion of asymptotical stability for index-1 tractable DAE (p.313)

Short Communications  Friday, 21 August
Wavelets III (organized by Andreas Rathsfeld)  H 2033

Chair: Pierre Gilles Lemarié-Rieusset

16:15 – 16:30 Siegfried Prößdorf, Weierstraß Institut, Berlin, Germany Boundary element methods using multiwavelets (p.310)

16:30 – 16:45 Jörg Schult, Weierstraß Institut, Berlin, Germany Stability condition for spline and multiwavelet collocation (p.312)

16:45 – 17:00 Marie Farge, École Normale Supérieure, Paris, France How to compute and model two-dimensional turbulent flows using wavelets?

Partial Differential Equations III  H 2033

Chair: Jianping Zhu

17:15 – 17:30 P. Sobolevskii, The Hebrew University of Jerusalem, Jerusalem, Israel The “counter-example” to W. Littman’s counter-example of L_{p}-energetical inequality for wave equation (p.314)

17:30 – 17:45 Ilja Schmelzer, WIAS, Berlin, Germany Grid Generation with Contravariant Geometry Description (p.311)

Short Communications  Saturday, 22 August
Partial Differential Equations IV  H 2033

Chair: Jennifer J. Zhao

14:00 – 14:15 Jennifer J. Zhao, University of Michigan, Dearborn, Michigan, USA Finite-difference approximations to a parabolic system modeling the thermoelastic contacts of two rods (p.319)

14:15 – 14:30 Jianping Zhu, Mississippi State University, Mississippi, MS, USA A High Order Algorithm for Solving Reaction Diffusion Equations (p.319)

14:30 – 14:45 Raul Kangro, Tartu University, Tartu, Estonia Analysis of Artificial Boundary Conditions for Black-Scholes Equations (p.304)

Partial Differential Equations V  H 2033

Chair: Cesar Palencia

15:00 – 15:15 Cesareo González, Universidad de Valladolid, Valladolid, Spain The implicit Euler method for abstract time-dependent parabolic problems with variable domains

(p.301)

15:15 – 15:30 Raimondas Ciegis, Institute of Mathematics and Informatics, Vilnius, Lithuania Odd — even difference schemes for the Schrödinger problem (p.298)

15:30 – 15:45 Daisuke Furihata, RIMS, Kyoto University, Kyoto, Japan Finite difference schemes that inherit the energy conservation or dissipation property from partial differential equations (p.300)

Short Communications  Monday, 24 August
Inverse Problems I  H 2033

Chair: Anatoly G. Yagola

14:00 – 14:15 Gottfried Bruckner, WIAS, Berlin, Germany On the regularization of an inverse vibration problem with noisy data (p.297)

14:15 – 14:30 Cesar Palencia, Universidad de Valladolid, Valladolid, Spain A numerical method for the integration of backward parabolic problems (p.308)

14:30 – 14:45 Sergii Pereverziev, Institute of Mathematics, Kiev, Ukraine An adaptive discretization of third kind Volterra equations with weakly singular kernels (p.309)

Inverse Problems II  H 2033

Chair: Sergii Pereverziev

15:00 – 15:15 Sergiy Georgievich Solodkyy, Ukrainian Acad. of Sci., Kiev, Ukraine Optimization of projection methods for solving linear ill-posed problems (p.314)

15:15 – 15:30 Anatoly G. Yagola, Moscow State University, Moscow, Russia Ill-posed problems and a priori information (p.318)

15:30 – 15:45 Weichung Wang, National Tainan Teachers College, Tainan, Taiwan Using Preconditioned Conjugate Gradients in Final Iterations of Interior Point Methods (p.317)

Short Communications  Tuesday, 25 August
Partial Differential Equations VI  H 2033

Chair: Ivanka P. Nikolova

16:15 – 16:30 Ivanka P. Nikolova, University of Pittsburgh, Pittsburgh, PA, USA Modeling of Fluid Flows with Strong Non-Linearity (p.307)

16:30 – 16:45 Christian von Törne, Universität Bonn, Bonn, Germany Method of Transport — A Multidimensional Scheme for Nonlinear Hyperbolic Systems (p.317)

16:45 – 17:00 Nicolai von Schroeders, UGH Paderborn, Paderborn, Germany Die numerische Lösung der schwachen Formulierung des Stokes-Systems im Außenraumgebiet

(p.317)

Partial Differential Equations VII  H 2033

Chair: Ursula van Rienen

17:15 – 17:30 Maxim A. Olshanskii, Moscow State University, Moscow, Russia Iterative method for generalized Stokes problem (p.308)

17:30 – 17:45 Chi-Wang Shu, Brown University, Providence,RI, USA High Order Weighted ENO Schemes for Unstructured Meshes (p.314)

17:45 – 18:00 Ursula van Rienen, Universität Rostock, Rostock, Germany Efficient Algorithms in Electromagnetic Field Computation (p.316)

Short Communications  Wednesday, 26 August
Wavelets IV (organized by Andreas Rathsfeld)  H 2033

Chair: Andreas Rathsfeld

16:15 – 16:30 Jürgen Prestin, GSF-Nat. Res. Center for Environment and Health, Neuherberg, Germany Polynomial wavelets

16:30 – 16:45 Andreas Rathsfeld, WIAS, Berlin, Germany Piecewise linear wavelet collocation for boundary element methods (p.310)

Poster Session  Saturday, 22 August 16:15 – 18:00  MA 144

Mansour Abdullah Al-Zanaidi, Kuwait University, Kuwait, Kuwait

A Class of Generalized Trapezoidal Formulas for the Integration of y’=f(x, y) (p.296) Mehmet Bozer, Eastern Mediterranean University, Famagusta, North Cyprus, Cyprus

The composite grids method of increased accuracy in solving the mixed boundary value problem for the Laplace equation (p.297)

Chin-Yun Chen, Universität Karlsruhe, Karlsruhe, Germany VANI: A package for Validated, Adaptive Numerical Integration in one and more dimensions (p.297)

Adigezal Dosiyev, Eastern Mediterranean University, Famagusta, North Cyprus, Cyprus A block-grid method for solving Laplace’s boundary value problem with singularities (p.298)

Bratislav Iricanin, University of Belgrade, New Belgrade, Yugoslavia

On the modification of Euler-Maclaurin summation formula connected with the Chebyshev quadrature formulae (p.302)

Anna Kamont, Mathematical Institute of the Polish Academy of Sciences, Sopot, Poland

Ditzian — Totik moduli of smoothness, spline systems and generalized Hölder spaces (p.303) Ivan G. Kazantsev, University Gent, Gent, Belgium

Tomographic reconstruction from arbitrary directions (p.304) Ingemar Kinnmark, Independent Researcher, Pittsburgh, PA, USA

Overdetermined Differential Equation Problems (p.304) Peeter Oja, Tartu University, Tartu, Estonia

Adaptive by monotonicity interpolating splines (p.307) Arvet Pedas, Tartu University, Tartu, Estonia

Collocation method for weakly singular integral equations (p.309) Gerd Reinhardt, Weierstraß-Institut, Berlin, Germany

Schnelle Algorithmen zum Clustern großer Datenmengen (p.311) Mayumi Shoji, Nihon University, Tokyo, Japan

Bifurcation structure of steady, deep-water gravity waves with constant vorticity (p.313) Svetlana N. Sytova, Institute of Nuclear Problems, Minsk, Republic of Belarus

Mathematical Modelling of X-ray Dynamical Diffraction by Non-Stationary Deformed Crystals (p.315) Zoltan Istvan Szabo, Kossuth University, Debrecen, Hungary

Root-finding Inclusion Algorithms Based on Tangential Lines and Parabolas (p.315) Svetlana B. Zaitseva, Moscow Power Engineering Institute, Moscow, Russia

On sharp error estimates in L_{2}-norm for methods of solving the heat equation in non-smooth data classes (p.318) Aihui Zhou, Academia Sinica, Beijing, China

Multiparameter error resolution for finite elements (p.319)